The perimeter of a rectangle 36m. The width is 18 yd less than twice the length. Find the length and width of the rectangle.



Answer by checkley77(12844)

W=2L
P=2W+2L
36=2(2L)+2L
36=4L+2L
36=6L
L=36/6
L=6 ANS. FOR THE LENGTH.
W=2*6=12 ANS. FOR THE WIDTH.
pROOF:
36=2*12+2^6
36=24+12
36=36